James Investes a part of 150,000 at 75 and the remaining part at 5%. How much did he invest at 7% if he receive a total interest of 10,000 after a year?
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52781)

.

Make correction.

Then re-post.



Answer by ankor@dixie-net.com(22740)

James Invests a part of 150,000 at 7% and the remaining part at 5%.
How much did he invest at 7% if he receive a total interest of 10,000 after a year? (Corrected)
:
let x = amt invested at 7%
The total is 150000, therefore
(150000-x) = amt invested at 5%
:
.07x + .05(150000-x) = 10000
.07x + 7500 - .05x = 10000
.07x - .05x = 10000 - 7500
.02x = 2500
x = 2500/.02
x = $125,000 invested at 7%
:
:
Check: 150000 - 125000 = 25000
.07(125000) = 8750
.05(25000) = 1250
-----------------------
total return 10000